# PsOCR: Benchmarking Large Multimodal Models for Optical Character Recognition in Low-resource Pashto Language

### v1

@@ 5.1 Models Performance Comparison
The evaluation of LMMs for Pashto OCR demonstrates substantial variability in performance across different metrics. A summary of the evaluation results is given in Table 2, accompanied by a graphical representation in Figure 9. Among the evaluated models, Gemini consistently achieved superior performance, obtaining the highest CA of 89.92%, WA of 69.5%, BLEU score of 49.54%, and an overall average of 69.65%. Conversely, the InternVL model exhibited the lowest performance, with an average score of -355.32%. Within the group of open-source models, Qwen-7B significantly outperformed the others, achieving a CA of 66.33%, WA of 27.41%, BLEU of 10.30%, and an average score of 34.68%. Among proprietary models, Grok showed the lowest performance, with an average of -16.66%. Notably, apart from Gemini, none of the proprietary models surpassed the open-source Qwen-7B in character accuracy, not even GPT-4o. Another important observation is that all models exhibited notably higher CA compared to WA and BLEU. This suggests that while these models can recognize individual characters, they struggle to generate coherent word sequences and phrases. The findings emphasize that Gemini holds considerable promise for applications demanding high-accuracy OCR without additional training. Meanwhile, for scenarios involving fine-tuning or further model customization, Qwen may be suitable due to its strong baseline performance and open-source availability.
Model CA (%) WA (%) BLEU (%) Avg (%)
InternVL -457.08 -608.94 0.06 -355.32
Janus -260.52 -261.86 0.03 -174.12
CPM -188.44 -193.65 0.84 -127.08
Llama -30.94 -98.74 5.64 -41.35
Florence-ft 6.29 -12.98 0.03 -2.22
Florence 11.94 -3.64 0.03 2.77
Qwen-3B 50.24 10.68 6.03 22.32
Qwen-7B 66.33 27.41 10.30 34.68
Grok -12.25 -38.16 0.44 -16.66
Claude 63.75 32.73 17.77 38.08
GPT-4o 65.92 35.53 27.01 42.82
Gemini 89.92 69.50 49.54 69.65
Table 2: Evaluation Results of LMMs on Pashto OCR Benchmark across different matrices
Figure 9: Performance comparison of evaluated LMMs on the Pashto OCR benchmark across different evaluation metrics
Figure 10: Effect of Image Size, Aspect Ratio, Font Size, Font Weight, Line Height and Text Alignment on Models’ Performance
Figure 11: Effect of Text Length, Theme, Colors and Font Families on Models’ Performance
Figure 12: Comparative Analysis of Models’ Scores for Top-10 Best-performing Font Families
Rank Font Name Avg Score Font Style
1. samsungonearabic-400 51.76 [Uncaptioned image]
2. niew-cromagnon 51.05 [Uncaptioned image]
3. ge-ss-text-medium 50.21 [Uncaptioned image]
4. mada 49.98 [Uncaptioned image]
5. calibrib 49.81 [Uncaptioned image]
6. abdomaster-normal 49.76 [Uncaptioned image]
7. vazir 49.70 [Uncaptioned image]
8. alghadtv 49.49 [Uncaptioned image]
9. psuedo-saudi 49.41 [Uncaptioned image]
10. droid-arabic-kufi 49.34 [Uncaptioned image]
Table 3: List of Top-10 Best-performing Font-families
@@ 5.2 In-depth Results Analysis:
To gain deeper insights into how various image and text properties influence Pashto OCR performance, we plotted each model’s average score against key image attributes, as shown in Figure 10. The following discussion summarizes the effect of each attribute on models’ performance.
@@
Effect of Font Family: Font family exerts the strongest influence on OCR accuracy, as shown in Figure 10 (J). Given the dataset’s diverse font families, models exhibit wide performance variation. These findings highlight font diversity as one of the primary challenges in Pashto OCR. Figure 11 presents a comparative analysis of models’ performance on the top 10 fonts, and Table 3 lists those fonts.
@@ 6 Conclusion and Future Work
In this work, we presented the first large-scale evaluation of eleven state-of-the-art LMMs on a newly developed Pashto OCR benchmark. Our dataset, comprising one million synthetically generated images annotated with bounding boxes at page, line, and token levels, represents the first publicly available resource of its kind for the low-resource Pashto language. We assessed seven open-source models (DeepSeek’s Janus, InternVL, MiniCPM, Florence, Qwen-7B, and Qwen-3B) alongside four proprietary models (GPT-4o, Gemini, Claude, and Grok) under zero-shot conditions. Experimental results revealed that Gemini delivered the highest overall performance, achieving 89.92% character accuracy, 69.5% word accuracy, and a BLEU score of 49.54%. Among open-source models, Qwen-7B stood out, significantly outperforming other freely available models. These findings underscore the zero-shot capabilities of current LMMs on cursive, complex scripts like Pashto and highlight the promise of open-source models such as Qwen for subsequent fine-tuning.
@@
Looking ahead, we are extending this work in two key directions. We are currently developing a Pashto VQA dataset and benchmark, and we are also working on the first large-scale handwritten Pashto OCR dataset. Together, these efforts will deepen our understanding of multimodal model performance on low-resource languages and drive future improvements in Pashto document analysis.

### v2

@@ 5.1 Models Performance Comparison
Tables 4 and 5, together with the visualization in Figure 8, present the performance of the evaluated LMMs on the PsOCR benchmark. To provide a comprehensive assessment, we report results using two complementary categories of evaluation metrics: transcription accuracy metrics, namely CER and WER, and text similarity metrics, including BLEU, METEOR, and BoW similarity. This dual evaluation framework enables analysis of both exact transcription correctness and overall textual similarity between predicted outputs and ground-truth documents.
The transcription accuracy results in Table 4 reveal substantial performance differences among the models. Gemini achieves the best overall performance, with a CER of 0.10 and a WER of 0.31, indicating strong character- and word-level transcription accuracy in a zero-shot setting. Among the proprietary models, GPT-4o and Claude also demonstrate competitive performance. Within the open-source group, Qwen-7B emerges as the strongest performer, achieving a CER of 0.34 and a WER of 0.73, closely matching the performance of some proprietary systems. In contrast, models such as Llama and Grok exhibit high error rates, with CER and WER values exceeding 1.0. Across all models, WER is consistently higher than CER, suggesting that while character recognition is relatively robust, maintaining correct word boundaries and producing accurate word sequences remains a significant challenge for Pashto OCR.
The text similarity results reported in Table 5 largely align with the transcription accuracy trends, while offering complementary insights into global textual similarity. Gemini again outperforms all other models, achieving the highest scores, with an average similarity of 0.63. GPT-4o and Claude follow, with GPT-4o showing stronger overall similarity to the ground truth, particularly in terms of lexical overlap. Among open-source models, Qwen-7B consistently attains the highest similarity scores, reinforcing its position as a strong open-source baseline. It is important to note that these similarity metrics provide auxiliary evidence of how closely the generated text resembles the ground truth at a sequence and lexical level.
These results highlight several key observations. First, proprietary LMMs generally outperform open-source alternatives; however, Qwen-7B significantly narrows this gap and, in some cases, approaches the performance of closed-source models. Second, the persistent disparity between CER and WER across models underscores the difficulty of word-level OCR for Pashto, a language characterized by cursive writing, ligatures, and ambiguous word boundaries. Third, the strong correspondence between low transcription error rates and high text similarity scores suggests that models producing accurate transcriptions also tend to generate text that is globally coherent and lexically aligned with the ground truth. Collectively, these findings demonstrate that Gemini offers the strongest zero-shot OCR performance for Pashto, while Qwen-7B represents a promising and accessible open-source alternative, establishing solid baselines for future research on OCR in low-resource, Perso-Arabic script languages.
Table 4: Comparative evaluation of LMMs on the Pashto OCR benchmark using CER and WER (lower is better).
      Model       CER       WER       Avg
      Llama       1.31       1.99       1.65
      Florence       0.88       1.04       0.96
      Qwen-3B       0.50       0.89       0.70
      Qwen-7B       0.34       0.73       0.53
      Grok       2.16       2.30       2.23
      Claude       0.36       0.67       0.52
      GPT-4o       0.30       0.60       0.45
      Gemini       0.10       0.31       0.20
Figure 8: Visualization of CER and WER performance of LMMs on the Pashto OCR benchmark (lower is better).
Table 5: Comparative evaluation of LMMs on the Pashto OCR benchmark using BLEU, METEOR, and BoW similarity metrics (higher is better).
    Model     BLEU     METEOR     BoW     Avg
    Llama     0.07     0.23     0.27     0.19
    Florence     0.00     0.00     0.00     0.00
    Qwen-3B     0.08     0.22     0.25     0.18
    Qwen-7B     0.12     0.32     0.34     0.26
    Grok     0.00     0.01     0.02     0.01
    Claude     0.19     0.36     0.37     0.31
    GPT-4o     0.28     0.49     0.49     0.42
    Gemini     0.50     0.72     0.67     0.63
Figure 9: Effect of different variations on models’ performance
Figure 10: Performance variation across fonts
Figure 11: LMMs scores on top-10 best-performing fonts
Table 6: List of top-10 best-performing font families
Rank Font Name Avg. CER Font Style
1. samsungonearabic-400 0.101 [Uncaptioned image]
2. vazir 0.109 [Uncaptioned image]
3. ge-$s$-text-medium 0.113 [Uncaptioned image]
4. frutiger-lt-arabic-45-light 0.115 [Uncaptioned image]
5. Vazirmatn-Regular 0.118 [Uncaptioned image]
6. boutrosnewsh1-bold 0.121 [Uncaptioned image]
7. sst-arabic-medium 0.126 [Uncaptioned image]
8. trumpit 0.126 [Uncaptioned image]
9. buka-puasa-bersama-normal 0.128 [Uncaptioned image]
10. Droid Naskh-Regular 0.130 [Uncaptioned image]
Figure 12: Example outputs of various models.
@@ 5.2 In-depth Results Analysis:
To gain deeper insights into how various image and text properties influence Pashto OCR performance, we plotted each model’s average score against key image attributes, as shown in Figure 9. Figure 12 lists some example outputs of various models. The following discussion summarizes the effect of each attribute on models’ performance.
@@
Effect of Font Family: Font family exerts the strongest influence on OCR accuracy, as shown in Figure 11. Given the dataset’s diverse font families, models exhibit wide performance variation. These findings highlight font diversity as one of the primary challenges in Pashto OCR. Figure 11 presents a comparative analysis of models’ performance on the top 10 fonts, and Table 6 lists those fonts.
@@ 6 Conclusion and Future Work
In this work, we presented the first large-scale evaluation of state-of-the-art LMMs on a newly developed Pashto OCR benchmark. Our dataset, comprising one million synthetically generated images annotated at the word, line, and document levels, represents the first publicly available resource of its kind for systematic OCR evaluation in the low-resource Pashto language. We evaluated four open-source models (Llama, Florence, Qwen-3B, and Qwen-7B) alongside four proprietary models (GPT-4o, Gemini, Claude, and Grok) under zero-shot settings using CER, WER, BLEU, METEOR, and BoW similarity metrics. Experimental results show that Gemini consistently achieves the best overall performance, attaining the lowest CER of 0.10 and WER of 0.31 and the highest text similarity scores, while among open-source models, Qwen-7B stands out as the strongest performer. These findings underscore both the growing zero-shot OCR capabilities of current LMMs for cursive, ligature-rich scripts such as Pashto and the promise of open-source models like Qwen-7B as strong foundations for future fine-tuning and adaptation.
@@
Looking ahead, we are extending this work in two key directions. First, we are currently developing a Pashto VQA dataset and benchmark. Second, we are working on the first large-scale handwritten Pashto OCR dataset. Furthermore, we plan to develop a more realistic version of the PsOCR dataset by adding different background patterns and diverse lighting conditions to the images, to more accurately mimic real-world conditions. Together, these efforts will deepen our understanding of multimodal model performance on low-resource languages and drive future improvements in Pashto document analysis.